OpenStudy (anonymous):

The width of a rectangle is 6 in. less than its length. The perimeter is 68 in. What is the width of the rectangle? in.

OpenStudy (anonymous):

I think its 14

Vocaloid (vocaloid):

nope, let's start from the beginning

Vocaloid (vocaloid):

let L = length let W = width W = L - 6, with me so far?

OpenStudy (anonymous):

yes

Vocaloid (vocaloid):

ok, now we use the formula for perimeter perimeter = 68 = 2(L+W), with me so far?

OpenStudy (anonymous):

yup:)

Vocaloid (vocaloid):

good, now we can rewrite our formula we know that W = L - 6, so we can replace "W" with "L-6" whenever we see it perimeter = 2(L+W) = 2(L+L-6), understand so far?

OpenStudy (anonymous):

Lol Ya

Vocaloid (vocaloid):

good, now let's put L + L together to get 2L perimeter = 2(2L-6) = 68 know how to solve for L now?

Vocaloid (vocaloid):

actually, you were right, W = 14, my bad ;_:

OpenStudy (anonymous):

Lol I thought so... I'll still give you a medal:) @Vocaloid